if the straight lines 3x-5y=7 and 4x+ay+9=0 are perpendicular to one another,find the value of a.

Grade:10

2 Answers

Arun
6 years ago
Dear student
 
If two lines are perpendicular then
Product of their slopes = -1
 
Hence
(3/5) *(-4/a) = -1
 
-12/5a = -1
a = 12/5
 
Regards
Arun (askIITians forum expert)
Latheef
6 years ago
Dear student If two lines are perpendicular thenProduct of their slopes = -1 Hence(3/5) *(-4/a) = -1 -12/5a = -1a = 12/5